# Training-Free Adaptation of New-Generation LLMs using Legacy Clinical Models

### v1

@@ 1 Introduction
Despite advances in general-domain language models, their application to clinical settings remains limited by hallucinations, omission of critical details, and failures in clinical reasoning (lehman2023dostillneed; hager2024llm_clinical_limits; asgari2025safety_llms). These shortcomings arise because pretraining corpora contain limited representations of clinical data, such as electronic health records, due to privacy constraints (singhal2023nature_clinical_knowledge).
Domain adaptation techniques such as continued pretraining and finetuning address these challenges but require resource-intensive retraining for each new model generation (e.g., MedPalm 1 $\rightarrow$ MedPalm 2) (singhal2025naturemed_mqa). This creates a lag between advances in base model capabilities and their clinical applicability, with substantial computational resources reinvested in domain adaptation for each new architecture.
We explore model ensembling as a training-free approach to combine the advanced reasoning of state-of-the-art general-domain models with the domain knowledge of legacy clinical models. Prior works show that integrating decoding-time probability distributions of general-domain models with heterogeneous architectures can yield efficient performance gains yao2025determinethenensemblenecessitytopkunion. However, these approaches assume largely overlapping model capabilities and have not been explored for combining models with distinct strengths, where information must be selectively integrated rather than uniformly aggregated. For example, we seek to incorporate the learned clinical knowledge of an older-generation clinical model without inheriting its limitations in instruction following, reasoning, or robustness due to its architecture or catastrophic forgetting (Kirkpatrick_2017).
Contrastive decoding offers a mechanism for isolating a model’s strengths by choosing tokens that maximize the likelihood difference between expert and amateur models, amplifying the expert’s behavior (li-etal-2023-contrastive). Proxy tuning leverages contrastive decoding to efficiently tune a large pretrained model by combining its logit distribution with the distributional delta between a smaller fine-tuned model and its untuned, base counterpart (liu2024proxy). However, proxy tuning assumes shared tokenization across models, restricting architectural diversity and preventing reuse of domain-adapted models built on older architectures.
@@ 1.1 Related Works
Model Ensembling. Probability-level model ensembling methods for heterogeneous architectures have primarily been explored with multiple general-domain models (xu2024bridginggapdifferentvocabularies; huang2024ensemblelearningheterogeneouslarge). UNiTE, the strongest-performing method to date, unions each model’s top-$k$ tokens at every decoding step and combines probabilities via re-tokenization (yao2025determinethenensemblenecessitytopkunion). Prior analyses show ensembling methods only outperform the ensemble’s strongest model when performance gaps are within 10% (yao2025determinethenensemblenecessitytopkunion). This limits applicability when combining new- and old-generation architectures where gaps can arise from architectural advances alone (bedi2025medhelm). In contrast, CAPT is designed for models with asymmetric capabilities and enables selective integration of learned information from older domain models.
Contrastive Decoding. Contrastive decoding has been used to steer pretrained models toward desirable behaviors, including reduced toxicity (liu-etal-2021-dexperts), increased helpfulness (mitchell2023eft), selective unlearning (suriyakumar2025ucdunlearningllmscontrastive), and improved coding, QA, and math performance (liu2024proxy). These methods require shared vocabularies, and are evaluated in task- or instruction-tuning settings that are in-distribution of the large model’s training corpora. Our approach supports models with disjoint vocabularies and is evaluated on clinical domain adaptation, a substantially out-of-distribution setting (Kim_2025).
@@ 2 Methods
which adds a correction representing the isolated clinical-domain signal. The selected token, $i^{*}=\arg\max_{i\in\mathcal{C}_{t}}s(i)$, is appended to the context. We set $k=20$, reflecting the top-20 log-probabilities exposed by black-box API providers (openai2025gpt5; google2025gemini3) and fix $\alpha=1.0$. See Appendix B for hyperparameter analysis.
Models and Baselines. We use Qwen3-30B (yang2025qwen3technicalreport) as the new-generation general-domain model ($M_{\text{new}}$) and MeLLaMA-13B-chat (xie2024mellama) as the old-generation clinical model ($M_{\text{old-clin}}$). MeLLaMA-13B-chat is continually pre-trained from LLaMA-2-13B-base ($M_{\text{old}}$) and instruction-tuned on clinical notes and tasks. We compare against proxy tuning (liu2024proxy) and UNiTE (yao2025determinethenensemblenecessitytopkunion), the strongest probability-level ensembling method to date. Since proxy tuning requires a shared tokenizer between models, we use LLaMA-2-70B-chat ($M_{\text{old-L}}$) (touvron2023llama2openfoundation) as its general-domain model.
@@
Evaluation Tasks. For classification, we report Macro-F1 and accuracy on 200 random test samples from: MedNLI (romanov2018lessonsnaturallanguageinference), classifying premise-hypothesis pairs; MTS-Specialty (MTSamples), identifying medical specialties from transcriptions; and CLIP (mullenbach2021clipdatasetextractingaction), extracting physician action items from discharge notes. For text generation, we report MedHELM LLM-jury scores (bedi2025medhelm) and MedVAL percent of risk-free outputs (aali2025medvalexpertlevelmedicaltext), both validated against clinician judgments, on 100 samples from: MIMIC-BHC (Aali_2024), generating brief hospital course summaries; MTS-Procedures (MTSamples), generating treatment plans; and MIMIC-RRS (Chen_2023), generating radiology study summaries. See Appendix C for further details on metrics and tasks.
@@ 5 Limitations
Evaluation. Evaluating the clinical utility of open-ended LLM-generated outputs remains difficult without manual expert review, which is costly and does not scale. While classification tasks provide standardized, widely used benchmarks, clinical LLMs are more often deployed in text-generation tasks (raji; EpicAIforClinicians2025; ShahLab2025ChatEHR; Small2024LLMInBasket), as smaller, simpler models can already achieve strong performance on basic classification tasks (li2023comparative). In our text-generation tasks, we report peer-reviewed evaluation metrics: MedHELM LLM-jury scores assess accuracy, clarity, and completeness against gold-standard references (bedi2025medhelm) and the percentage of risk-free outputs as determined by MedVAL, a fine-tuned clinical evaluation model (aali2025medvalexpertlevelmedicaltext). While both metrics have been benchmarked against physician reviews, they provide limited insight into the multi-dimensional nature of output quality. In practice, we observed that the MedHELM LLM jury reasoning for each score differed substantially from physician analysis, as the LLM jury did not highlight specific contextual accuracies or information.
@@ 6 Ethical Considerations
Our tasks are derived from two clinical datasets: MIMIC (johnson2016mimic) and MTSamples (MTSamples). MIMIC is a freely available, de-identified dataset of patients admitted to critical care units at Beth Israel Deaconess Medical Center. MTSamples is a publicly available, de-identified collection of medical transcription sample reports contributed by users for reference purposes. In terms of data consent, MIMIC data was collected under its respective IRB approval where the requirement for individual patient consent was waived. We access MIMIC under a data use agreement. MTSamples consists of voluntarily contributed data. Neither dataset requires Ethics Review Board approval for secondary analysis, as we perform in this study. We do not collect any additional data for this study which would require Ethics Review Board approval.
@@ Appendix A Additional Related Work
Clinical Domain Adaptation. Clinical models such as Gatortron, clinicalBERT, NYUTron, Me-LLaMA, and Clinical-T5 have demonstrated the effectiveness of continued pretraining on large volumes of unlabeled medical text (yang2022gatortronlargeclinicallanguage; alsentzer2019publiclyavailableclinicalbert; jiang2023healthsystem; xie2024mellama; lehman2023clinicalt5). Following continued pretraining, supervised fine-tuning can further improve performance on specific clinical tasks, instruction following, and alignment with human preferences (han2023medalpaca; zhang2023alpacare; singhal2023nature_clinical_knowledge; wang2024clinicnote; zhang2023huatuogpt). While effective, these approaches are highly resource-intensive, especially continued pre-training: for example, Me-LLaMA conducted continued pretraining on LLaMA-2 base models using 129B tokens and 160 × 80 GB A100 GPUs. Our work complements these techniques by addressing their limited transferability, as the benefits of such costly adaptations are typically confined to the specific model on which they are performed.
@@ Appendix B Hyperparameter Analysis
To evaluate the effect of $k$ and $\alpha$, we conducted experiments on two held-out datasets. We report Macro-F1 on the validation split of the MedNLI classification task, and MedHelm LLM-jury score on ACI-Bench, a text-generation task of generating a structured clinical note from patient-doctor dialogues (yim2023acibenchnovelambientclinical). We use ACI-Bench because MedHelm does not provide a validation set for its text-generation tasks and we did not want to reduce sample size of our tasks. Table 2 shows that varying $k$ has minimal impact on performance. We select $k=20$ because black-box API providers provide the top-20 logprobs for each token generation (openai2025gpt5; google2025gemini3). We set $\alpha=1.0$ because it achieves best performance on both tasks, as shown by Table 3.
@@ Appendix C Experimental Setup
For text-generation tasks, we implement a custom model client within the MedHELM evaluation framework (bedi2025medhelm). We use MedHELM because it provides a standardized, reproducible, and peer-reviewed evaluation pipeline across clinical tasks, and includes an LLM-jury metric for text-generation tasks which assesses on medical accuracy, completeness and clarity compared to the reference output. Through use of MedHELM, our results can be directly compared to public leaderboard results, and CAPT can be readily applied to any of the 37 MedHELM clinical classification, text generation, or question-answering tasks. In addition, we report the percentage of responses that received a Level 1 (risk-free) grade using MedVAL, a physician-aligned risk grading LLM-as-a-judge system (levels 1–4, where level 1 is risk-free and level 4 represents the highest level of clinical risk). This metric assesses the clinical utility of text-generation outputs at scale, in addition to our qualitative analyses.
@@ Appendix E Physician Case Study
We present five randomly selected full-length CAPT outputs from the MTS-Procedure task, annotated with token-level shifts and expert qualitative analysis. The analysis was conducted by two authors of this paper, an interventional radiologist with 25 years of experience writing post-operative management plans and an obstetric gynecologist with expertise in clinical language models. Both were provided verbal instructions. We use MTS-Procedure because it is publicly available; our other text-generation tasks rely on MIMIC (johnson2016mimic), whose raw data cannot be shared. Figures 471013, and 16 contain the annotated CAPT outputs with expert qualitative analysis. For each case, we provide the LLM jury evaluation (Figures 58111417), and the corresponding output from the new-generation general-domain model ($M_{\text{new}}$) (Figures 691215, 18).

### v2

@@ 1 Introduction
Despite advances in general-domain language models, their application to clinical settings remains limited by hallucinations, omission of critical details, and failures in clinical reasoning (Lehman et al., 2023; Hager et al., 2024; Asgari et al., 2025). These shortcomings arise because pretraining corpora contain limited representations of clinical data, such as electronic health records, due to privacy constraints (Singhal et al., 2023a).
Domain adaptation techniques such as continued pretraining and finetuning address these challenges but require resource-intensive retraining for each new model generation (e.g., MedPalm 1 $\rightarrow$ MedPalm 2) (Singhal et al., 2023b). This creates a lag between advances in base model capabilities and their clinical applicability, with substantial computational resources reinvested in domain adaptation for each new architecture.
We explore model ensembling as a training-free approach to combine the advanced reasoning of state-of-the-art general-domain models with the domain knowledge of legacy clinical models. Prior works show that integrating decoding-time probability distributions of general-domain models with heterogeneous architectures can yield efficient performance gains Yao et al. (2025). However, these approaches assume largely overlapping model capabilities and have not been explored for combining models with distinct strengths, where information must be selectively integrated rather than uniformly aggregated. For example, we seek to incorporate the learned clinical knowledge of an older-generation clinical model without inheriting its limitations in instruction following, reasoning, or robustness due to its architecture or catastrophic forgetting (Kirkpatrick et al., 2017).
Contrastive decoding offers a mechanism for isolating a model’s strengths by choosing tokens that maximize the likelihood difference between expert and amateur models, amplifying the expert’s behavior (Li et al., 2023a). Proxy tuning leverages contrastive decoding to efficiently tune a large pretrained model by combining its logit distribution with the distributional delta between a smaller fine-tuned model and its untuned, base counterpart (Liu et al., 2024). However, proxy tuning assumes shared tokenization across models, restricting architectural diversity and preventing reuse of domain-adapted models built on older architectures.
@@ 1.1 Related Works
Model Ensembling. Probability-level model ensembling methods for heterogeneous architectures have primarily been explored with multiple general-domain models (Xu et al., 2024; Huang et al., 2024). UNiTE, the strongest-performing method to date, unions each model’s top-$k$ tokens at every decoding step and combines probabilities via re-tokenization (Yao et al., 2025). Prior analyses show ensembling methods only outperform the ensemble’s strongest model when performance gaps are within 10% (Yao et al., 2025). This limits applicability when combining new- and old-generation architectures where gaps can arise from architectural advances alone (Bedi et al., 2025). In contrast, CAPT is designed for models with asymmetric capabilities and enables selective integration of learned information from older domain models.
Contrastive Decoding. Contrastive decoding has been used to steer pretrained models toward desirable behaviors, including reduced toxicity (Liu et al., 2021), increased helpfulness (Mitchell et al., 2023), selective unlearning (Suriyakumar et al., 2025), and improved coding, QA, and math performance (Liu et al., 2024). These methods require shared vocabularies, and are evaluated in task- or instruction-tuning settings that are in-distribution of the large model’s training corpora. Our approach supports models with disjoint vocabularies and is evaluated on clinical domain adaptation, a substantially out-of-distribution setting (Kim et al., 2025).
@@ 2 Methods
which adds a correction representing the isolated clinical-domain signal. The selected token, $i^{*}=\arg\max_{i\in\mathcal{C}_{t}}s(i)$, is appended to the context. We set $k=20$, reflecting the top-20 log-probabilities exposed by black-box API providers (OpenAI, 2025; Google DeepMind, 2025) and fix $\alpha=1.0$. See Appendix B for hyperparameter analysis.
Models and Baselines. We use Qwen3-30B (Yang et al., 2025) as the new-generation general-domain model ($M_{\text{new}}$) and MeLLaMA-13B-chat (Xie et al., 2024) as the old-generation clinical model ($M_{\text{old-clin}}$). MeLLaMA-13B-chat is continually pre-trained from LLaMA-2-13B-base ($M_{\text{old}}$) and instruction-tuned on clinical notes and tasks. We compare against proxy tuning (Liu et al., 2024) and UNiTE (Yao et al., 2025), the strongest probability-level ensembling method to date. Since proxy tuning requires a shared tokenizer between models, we use LLaMA-2-70B-chat ($M_{\text{old-L}}$) (Touvron et al., 2023) as its general-domain model.
@@
Evaluation Tasks. For classification, we report Macro-F1 and accuracy on 200 random test samples from: MedNLI (Romanov and Shivade, 2018), classifying premise-hypothesis pairs; MTS-Specialty (MTSamples, 2025), identifying medical specialties from transcriptions; and CLIP (Mullenbach et al., 2021), extracting physician action items from discharge notes. For text generation, we report MedHELM LLM-jury scores (Bedi et al., 2025) and MedVAL percent of risk-free outputs (Aali et al., 2025), both validated against clinician judgments, on 100 samples from: MIMIC-BHC (Aali et al., 2024), generating brief hospital course summaries; MTS-Procedures (MTSamples, 2025), generating treatment plans; and MIMIC-RRS (Chen et al., 2023), generating radiology study summaries. See Appendix C for details on metrics and tasks.
@@ 5 Limitations
Evaluation. Evaluating the clinical utility of open-ended LLM-generated outputs remains difficult without manual expert review, which is costly and does not scale. While classification tasks provide standardized, widely used benchmarks, clinical LLMs are more often deployed in text-generation tasks (Raji et al., 2025; Corporation, 2025; Shah Lab, 2025; Small et al., 2024), as smaller, simpler models can already achieve strong performance on basic classification tasks (Li et al., 2023b). In our text-generation tasks, we report peer-reviewed evaluation metrics: MedHELM LLM-jury scores assess accuracy, clarity, and completeness against gold-standard references (Bedi et al., 2025) and the percentage of risk-free outputs as determined by MedVAL, a fine-tuned clinical evaluation model (Aali et al., 2025). While both metrics have been benchmarked against physician reviews, they provide limited insight into the multi-dimensional nature of output quality. In practice, we observed that the MedHELM LLM jury reasoning for each score differed substantially from physician analysis, as the LLM jury did not highlight specific contextual accuracies or information.
@@ 6 Ethical Considerations
Our tasks are derived from two clinical datasets: MIMIC (Johnson et al., 2016) and MTSamples (MTSamples, 2025). MIMIC is a freely available, de-identified dataset of patients admitted to critical care units at Beth Israel Deaconess Medical Center. MTSamples is a publicly available, de-identified collection of medical transcription sample reports contributed by users for reference purposes. In terms of data consent, MIMIC data was collected under its respective IRB approval where the requirement for individual patient consent was waived. We access MIMIC under a data use agreement. MTSamples consists of voluntarily contributed data. Neither dataset requires Ethics Review Board approval for secondary analysis, as we perform in this study. We do not collect any additional data for this study which would require Ethics Review Board approval.
@@ Appendix A Additional Related Work
Clinical Domain Adaptation. Clinical models such as Gatortron, clinicalBERT, NYUTron, Me-LLaMA, and Clinical-T5 have demonstrated the effectiveness of continued pretraining on large volumes of unlabeled medical text (Yang et al., 2022; Alsentzer et al., 2019; Jiang et al., 2023; Xie et al., 2024; Lehman and Johnson, 2023). Following continued pretraining, supervised fine-tuning can further improve performance on specific clinical tasks, instruction following, and alignment with human preferences (Han et al., 2025; Zhang et al., 2025; Singhal et al., 2023a; Wang et al., 2025; Zhang et al., 2023). While effective, these approaches are highly resource-intensive, especially continued pre-training: for example, Me-LLaMA conducted continued pretraining on LLaMA-2 base models using 129B tokens and 160 × 80 GB A100 GPUs. Our work complements these techniques by addressing their limited transferability, as the benefits of such costly adaptations are typically confined to the specific model on which they are performed.
@@ Appendix B Hyperparameter Analysis
To evaluate the effect of $k$ and $\alpha$, we conducted experiments on two held-out datasets. We report Macro-F1 on the validation split of the MedNLI classification task, and MedHelm LLM-jury score on ACI-Bench, a text-generation task of generating a structured clinical note from patient-doctor dialogues (Yim et al., 2023). We use ACI-Bench because MedHelm does not provide a validation set for its text-generation tasks and we did not want to reduce sample size of our tasks. Table 2 shows that varying $k$ has minimal impact on performance. We select $k=20$ because black-box API providers provide the top-20 logprobs for each token generation (OpenAI, 2025; Google DeepMind, 2025). We set $\alpha=1.0$ because it achieves best performance on both tasks, as shown by Table 3.
@@ Appendix C Experimental Setup
For text-generation tasks, we implement a custom model client within the MedHELM evaluation framework (Bedi et al., 2025). We use MedHELM because it provides a standardized, reproducible, and peer-reviewed evaluation pipeline across clinical tasks, and includes an LLM-jury metric for text-generation tasks which assesses on medical accuracy, completeness and clarity compared to the reference output. Through use of MedHELM, our results can be directly compared to public leaderboard results, and CAPT can be readily applied to any of the 37 MedHELM clinical classification, text generation, or question-answering tasks. In addition, we report the percentage of responses that received a Level 1 (risk-free) grade using MedVAL, a physician-aligned risk grading LLM-as-a-judge system (levels 1–4, where level 1 is risk-free and level 4 represents the highest level of clinical risk). This metric assesses the clinical utility of text-generation outputs at scale, in addition to our qualitative analyses.
@@ Appendix E Physician Case Study
We present five randomly selected full-length CAPT outputs from the MTS-Procedure task, annotated with token-level shifts and expert qualitative analysis. The analysis was conducted by two authors of this paper, an interventional radiologist with 25 years of experience writing post-operative management plans and an obstetric gynecologist with expertise in clinical language models. Both were provided verbal instructions. We use MTS-Procedure because it is publicly available; our other text-generation tasks rely on MIMIC (Johnson et al., 2016), whose raw data cannot be shared. Figures 471013, and 16 contain the annotated CAPT outputs with expert qualitative analysis. For each case, we provide the LLM jury evaluation (Figures 58111417), and the corresponding output from the new-generation general-domain model ($M_{\text{new}}$) (Figures 691215, 18).